# Cost-Effective Online Multi-LLM Selection with Versatile Reward Models

### v1

@@ 4.2 Procedures by Scheduling Cloud
Discretization Rounding for LLM Selection. In our architecture, the scheduling cloud can communicate with each local server, while the local servers do not directly communicate with each other. As outlined in Section 3, our focus is on elucidating the one-to-one relationship between a local server and the cloud. With the fully original user data stored locally, the local server sends the relaxed continuous data $\tilde{\boldsymbol{Z}}_{t}$ to the scheduling cloud, which then coordinates various LLMs and selects a new action in response to the requests. Inspired by the works of [10, 22], the synchronized cloud employs specialized discretization rounding algorithms based on the reward models to convert the relaxed, continuous data back into the discrete form. By discretizing $\tilde{\boldsymbol{Z}}_{t}$, the scheduling cloud identifies a feasible set of LLMs, $S_{t}$, to schedule and select for the current round $t$ (line 7 in Algorithm 1). Details on the discretization rounding algorithms are provided below, tailored to the specific reward models.
@@
For the AWC reward, as detailed in Algorithm 2, the scheduling cloud initiates the process by setting up a vector $\boldsymbol{v}$ and identifying $K$ sets $\{\mathcal{B}_{1},\cdots,\mathcal{B}_{K}\}$ (line 2). Following this initialization, Algorithm 2 enters an iterative phase where it fine-tunes the composition of the sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$, aiming to keep the evolving solution within acceptable bounds while also seeking to optimize the reward (lines 4-10). Subsequent steps involve refining $\mathcal{B}_{1}$ by excluding all elements found in set $\mathcal{G}$, and updating $\mathcal{A}_{i+1}$ to reflect the current state of $\mathcal{B}_{1}$ (line 11), which are designed to gradually construct the final integrated solution. Finally, Algorithm 2 designates $\mathcal{A}_{K}$ as the final action of selected LLM $S_{t}$ (line 13). The validity of the resultant solution for LLM selection is underpinned by the lemma that follows (more discussions are in Section A.1).
@@
5 Performance Analysis
In this section, we conduct a comprehensive analysis of the theoretical performance (regret in Eq. (2), and violation in Eq. (1)) of C2MAB-V. Due to the page limit, we put the proof in Appendix B.
To state the analysis, we firstly give some definitions. Let $o_{t,S}$ represent the probability that all LLMs within a selected action $S$ are observed at round $t$, and $r^{*}=\max_{S\in\mathcal{S}}r(S;\boldsymbol{\mu})\leq NL$ be the maximum reward for the $L$-Lipschitz reward function $r(S;\boldsymbol{\mu})$. Following [32, 36], we define $o^{*}=\min_{t\in\mathcal{T},S\in\mathcal{S}}o_{t,S}$ as the minimum observation probability across all feasible combinations of LLM selection, within the context of partial feedback mechanisms.
Theorem 1 (Regret Bound).
$R(T)\leq\frac{4L}{o^{*}}\sqrt{2NKT\ln\left(\frac{2\pi^{2}KT}{3}\right)}+\left(K+1\right)r^{*}.$ (8)
Remark 1.
The C2MAB-V framework extends the capabilities of existing CMAB models by incorporating long-term cost considerations, which is typically not present in previous works [13, 34, 33]. When comparing with the linear CMAB model analyzed by [34], which focuses on full rather than partial feedback, by setting $L=1$ and $o^{*}=1$, our regret bound aligns with theirs, adhering to the lower bound $\Omega(\sqrt{NKT})$ in [34], up to a $\sqrt{\ln T}$ factor. Similarly, our findings are consistent with the regret bounds in studies of CMAB with combinatorial feedback [33, 32]. The comparability of our regret bound with other CMAB frameworks primarily stems from the efficient management of uncertainty in parameter estimation. A unique difference in our approach, however, is the implementation of a discretization process for the relaxed NP-hard problem. Despite this modification, due to our efforts on the shared attributes of different reward functions, the regret of C2MAB-V under the long-term cost constraint maintains the same order as those observed in other CMAB studies.
Theorem 2 (Violation Bound).
With $\delta=1/T$ in the confidence radius, the constraint violation within the C2MAB-V framework is bounded as follows, with probability at least $1-1/T$:
$V(T)\leq\sqrt{NK/T}\left(4\sqrt{2\ln\left(\frac{2\pi^{2}KT}{3}\right)}+\sqrt{SK/T}\right).$ (9)
Remark 2.
As discussed in Section 3, achieving a violation that diminishes rapidly is paramount. Our analysis reveals that the violation decreases at a rate of $\tilde{O}(\sqrt{\frac{K}{T}})$. As $T$ grows large, $V(T)$ approaches zero, suggesting an eventual elimination of violation. Furthermore, the overall violation is shown to be $\tilde{O}(\sqrt{KT})$, which is comparable to the order of regret. An interesting comparison arises with the work of [59], which studies non-linear submodular rewards and linear costs within the context of knapsack constraints, specifically focusing on scenarios with known and fixed costs. This contrasts with our exploration of scenarios characterized by unknown stochastic costs. Despite the inherent challenges posed by unknown costs, our framework manages to attain an $\tilde{O}(\sqrt{T})$ approximate regret with an approximation ratio of $\alpha_{1}=1-\frac{1}{e}\approx 0.632$ under knapsack constraints. In contrast, the method from [59] achieves an $\tilde{O}(\sqrt{T})$ approximate regret with a lower approximation ratio $\alpha_{2}\leq 0.5$. Thus, our approach secures a regret improvement of at least $\left(\alpha_{1}-\alpha_{2}\right)T\geq 0.132T$. Compared to [55], which focuses on addressing constraints, our work expands to include non-linear rewards. For linear cases, however, our approach closely mirrors that of [55].
6 Performance Evaluation
LLM_ID Model Name (with Parameters) Cost (USD/1k tokens) Size (GB)
1 ChatGLM2-6B-32K [47] 0.005 12.5
2 ChatGPT-3.5 [47] 0.02 /
3 Claude 2 [14] 0.08 /
4 ERNIE 3.5-8K [4] 0.015 /
5 Llama 2-7B [46] 0.005 12.6
6 Llama 2-13B [46] 0.008 24.3
7 Llama 2-70B [46] 0.05 128.3
8 Mixtral-8x7B-Instruct [1] 0.05 93.37
9 ChatGPT-4 [47] 0.12 /
Experiment Settings. We evaluate three multi-LLM reward models (AWC, SUC, AIC) to represent different task types. All experiments are conducted on a device equipped with an Intel Core i5-13600KF CPU @ 3.50GHz and 32 GB of memory. Gurobi, a powerful mathematical optimization solver supporting various programming models, is utilized to solve relaxed constraint optimization problems using version 11.0.1 [48]. Fig. 4 illustrates a sample interaction involving large language models (LLMs) across various scientific disciplines—Biology, Chemistry, Geography, and Physics—utilizing the SciQ dataset [66]. Table 2 lists the LLMs used, detailing each model’s name, parameters, cost per 1,000 tokens in USD, and size in gigabytes. The models range from smaller configurations, such as ChatGLM2-6B-32K, to larger ones like Llama 2-70B. Given the variability in generating answers when using LLMs and potential non-responses, we specify reward allocations based on different LLM return scenarios, comparing them with the SciQ dataset’s labels. If the question is answered accurately and adheres to the prescribed prompt format, the reward is set at 0.5. For incorrect responses, no reward is granted. However, if network congestion or LLM issues result in an empty response, a reward of 0.1 is allocated. If the response adheres to the format guidelines but fails to meet the specified requirements, a reward of 0.3 is assigned. Results are averaged over 10 seeds with a 95% confidence interval. The maximum number $N$ is uniformly set to 4, based on the LLM set size. Budget thresholds $\rho$ are 0.45 for AWC, 0.5 for SUC, and 0.3 for AIC, according to reward models and official LLM pricing. More comprehensive experiments (e.g., varying budget thresholds, impacts of maximum number, and comparisons with and without relaxation) are available in Appendix C.
Comparison Benchmarks. Comparisons include consistently utilizing the expensive ChatGPT-4 [47] and the cheap ChatGLM2 [47], the online CMAB algorithm CUCB, [63], which ignores constraints, and $\epsilon$-Greedy [2], which alternates between using empirical means and selecting uniformly based on the adaptive $\epsilon_{t}=\text{min}\{1,\frac{2\sqrt{K}}{\sqrt{t}}\}$ (Here, [55, 59] are not compared due to being restricted to fixed costs or linear reward). The robustness of C2MAB-V is validated by varying the parameters $\alpha_{\mu},\alpha_{c}$ with values of (0.3, 0.05), (1, 0.05), (0.3, 0.01), and (1, 0.01), referred to as (a), (b), (c), (d).
Performance Metric. To balance both reward and cost considerations, we assess performance using three metrics: reward, violation, and the reward/violation ratio. The reward/violation ratio is defined as the average per-round reward divided by the average per-round violation, mathematically represented as $\frac{\sum_{\tau=1}^{t}r(S_{\tau},\boldsymbol{\mu})/t}{\sum_{\tau=1}^{t}V(\tau)/t}$. Higher ratios indicate superior performance. Detailed experimental results, including variations in the budget threshold and a stronger focus on performance, are provided in Appendix C.1.
(a) AWC
(b) SUC
(c) AIC
Evaluation Results. Fig. 6 and Fig. 7 display the experimental rewards and violations for three different task types. In Fig. 6, C2MAB-V consistently outperforms or matches the $\epsilon$-Greedy baseline in terms of per-round rewards. The parameters on $\alpha_{\mu}$ and $\alpha_{c}$ used are (0.3, 0.05), (1, 0.05), (0.3, 0.01), and (1, 0.01), denoted respectively as (a), (b), (c), and (d). Adjusting the values of $\alpha_{\mu}$ and $\alpha_{c}$ results in consistently higher rewards for C2MAB-V settings. These settings also significantly reduce per-round violations compared to the $\epsilon$-Greedy approach. While the CUCB algorithm achieves the highest per-round rewards, it does so by completely ignoring cost constraints, leading to a per-round violation rate at least four times higher than that of C2MAB-V, as shown in Fig. 7. This figure also indicates that C2MAB-V’s violations are lower than those of the $\epsilon$-Greedy baseline, highlighting its ability to effectively manage constraints. Additionally, as shown in Fig. 6(c), we can observe a convergence trend in the rewards of C2MAB-V. This trend suggests that the initial actions selected might have had higher rewards but also higher violations. Over time, the cost-aware online learning strategy adjusts by slightly reducing rewards to achieve greater cost savings, demonstrating its effectiveness in managing long-term constraints.
The performance of the reward/ violation ratio is illustrated in Fig. 5. ChatGLM2 is excluded since its rewards are significantly low, below 0.18, 0.10, and 0.0001 in the AWC, SUC, AIC models, despite no violations. In the AWC model (Fig. 5(a)), C2MAB-V consistently outperforms other algorithms in all four parameter settings, achieving the highest reward/violation ratio. This underscores C2MAB-V’s superior ability to balance rewards against violations. The patterns observed for the SUC and AIC models (Fig. 5(b) and Fig. 5(c)) mirror those of the AWC model, further showing the robustness of C2MAB-V across different settings. For instance, when randomly evaluating C2MAB-V(c), it improves by at least 64.72% over $\epsilon$-Greedy, by a factor of four over CUCB, and by a factor of six over Always using ChatGPT-4 across three tasks. In summary, C2MAB-V maintains the highest reward/violation ratio under different $(\alpha_{\mu},\alpha_{c})$ for multi-LLM tasks, aligning with our theoretical analysis.
(a) AWC
(b) SUC
(c) AIC
Figure 6: Reward of three task types with nine different LLMs.
(a) AWC
(b) SUC
(c) AIC
Figure 7: Violation of three task types with nine different LLMs.
7 Conclusion
In this paper, we introduce C2MAB-V, a cost-effective combinatorial online model with a versatile reward structure, designed to efficiently select multiple LLMs based on specific task requirements while adhering to budget constraints. C2MAB-V incorporates continual online feedback, and transform the NP-hard multi-LLM selection problem into a manageable relaxed form, with a discretization rounding scheme utilized to coordinate multiple LLMs within a local-cloud architecture. Theoretical analysis of C2MAB-V provides robust guarantees for the framework’s efficacy, including sublinear regret and rapidly diminishing violation. Moreover, empirical evaluations with nine LLMs demonstrate C2MAB-V’s capability to balance performance with cost efficiency across three distinct types of collaborative tasks. Looking ahead, detailed aspects on privacy protection and communication between multiple local servers could be further developed within our local-cloud architecture.
Appendix A Constraint and Reward Analysis
A.1 Discussions on Constraint Type

### v2

@@ 4.2 Procedures by Scheduling Cloud
Discretization Rounding for LLM Selection. In our architecture, the scheduling cloud can communicate with each local server, while the local servers do not directly communicate with each other. As described in Section 3, our focus is on elucidating the one-to-one relationship between a local server and the cloud. With fully original user data stored locally, the local server sends the relaxed continuous data $\tilde{\boldsymbol{Z}}_{t}$ to the scheduling cloud, which then coordinates various LLMs and selects a new action in response to requests. Inspired by the works of [11, 24], the synchronized cloud employs specialized discretization rounding algorithms based on reward models to convert the relaxed continuous data back into discrete form. By discretizing $\tilde{\boldsymbol{Z}}_{t}$, the scheduling cloud identifies a feasible set of LLMs $S_{t}$ to schedule and select for the current round $t$ (line 7 in Algorithm 1). The corresponding algorithms are tailored solutions for our versatile reward model and represent a significant effort to find suitable methods for our specific setup. However, since the design of these algorithms is not our core contribution, details are provided in Appendix B. Note that discretization rounding can reduce the complexity of LLM combination selection, as detailed in the computational efficiency comparison with direct discrete constrained optimization in Appendix E.3.
5 Performance Analysis
In this section, we conduct a comprehensive analysis of the theoretical performance (regret in Eq. (2), and violation in Eq. (1)) of C2MAB-V. Due to the page limit, we put the proof in Appendix D.
To state the analysis, we firstly give some definitions. Let $o_{t,S}$ represent the probability that all LLMs within a selected action $S$ are observed at round $t$, and $r^{*}=\max_{S\in\mathcal{S}}r(S;\boldsymbol{\mu})\leq NL$ be the maximum reward for the $L$-Lipschitz reward function $r(S;\boldsymbol{\mu})$. Following [36, 40], we define $o^{*}=\min_{t\in\mathcal{T},S\in\mathcal{S}}o_{t,S}$ as the minimum observation probability across all feasible combinations of LLM selection, within the context of partial feedback mechanisms. For the AWC scenario, as previously discussed, a subset may be chosen from multiple available options for an action $S_{t}$. In such cases, we provide the violation upper bound under the worst-case scenario, where each LLM in action $S_{t}$ fails to satisfy the user until the last one, i.e., $S_{t}=\mathcal{F}_{t}$ for all $t\in\mathcal{T}$.
Theorem 1 (Regret Bound).
$R(T)\leq\frac{2L}{o^{*}}\sqrt{2NKT\ln\left(\frac{2\pi^{2}KT}{3}\right)}+\left(K+1\right)r^{*}.$ (7)
Remark 1.
The C2MAB-V framework extends the capabilities of existing CMAB models by incorporating long-term cost considerations, which is typically not present in previous works [14, 38, 37, 36]. When comparing with the linear CMAB model analyzed by [38], which focuses on full rather than partial feedback, by setting $L=1$ and $o^{*}=1$, our regret bound aligns with theirs, adhering to the lower bound $\Omega(\sqrt{NKT})$ in [38], up to a $\sqrt{\ln T}$ factor. The comparability of our regret bound with other CMAB frameworks primarily stems from the efficient management of uncertainty in parameter estimation. A unique difference in our approach, however, is the implementation of a discretization process for the relaxed NP-hard problem. Despite this modification, due to our efforts on the shared attributes of different reward functions (which is not an easy task), the regret of C2MAB-V under the long-term cost constraint maintains the same order as those observed in other CMAB studies.
Theorem 2 (Violation Bound).
With $\delta=1/T$ in the confidence radius, the constraint violation under the worst case is bounded as follows, with probability at least $1-1/T$:
$V(T)\leq\sqrt{NK/T}\left(2\sqrt{2\ln\left(\frac{2\pi^{2}KT}{3}\right)}+\sqrt{NK/T}\right).$ (8)
Remark 2.
As discussed in Section 3, achieving a violation that diminishes rapidly is paramount. Our analysis reveals that the violation decreases at a rate of $\tilde{O}(\sqrt{\frac{K}{T}})$. As $T$ grows large, $V(T)$ approaches zero, suggesting an eventual elimination of violation. Furthermore, the overall violation is shown to be $\tilde{O}(\sqrt{KT})$, which is comparable to the order of regret. An interesting comparison arises with the work of [64], which studies non-linear submodular rewards and linear costs within the context of knapsack constraints, specifically focusing on scenarios with known and fixed costs. This contrasts with our exploration of scenarios characterized by unknown stochastic costs. Despite the inherent challenges posed by unknown costs, our framework manages to attain an $\tilde{O}(\sqrt{T})$ approximate regret with an approximation ratio of $\alpha_{1}=1-\frac{1}{e}\approx 0.632$ under knapsack constraints. In contrast, the method from [64] achieves an $\tilde{O}(\sqrt{T})$ approximate regret with a lower approximation ratio $\alpha_{2}\leq 0.5$. Thus, our approach secures a regret improvement of at least $\left(\alpha_{1}-\alpha_{2}\right)T\geq 0.132T$. Compared to [59], which focuses on addressing constraints, our work expands to include non-linear rewards.
6 Performance Evaluation
Experiment Settings. We evaluate the three multi-LLM reward models (AWC, SUC, AIC) to represent different task types using the SciQ dataset [69] across nine LLMs: ChatGPT-3.5, ERNIE 3.5, ChatGPT-4, ChatGLM2, Llama 2-7B, Llama 2-13B, Llama 2-70B, Mixtral-8x7B, and Claude 2 [52, 15, 50, 5, 51]. [^3] Our evaluation focuses exclusively on online learning from continuous feedback, incorporating offline pre-training, requiring offline training or GPU resources, or addressing cold-start issues. Studies, such as [72], on accelerating online learning to overcome cold-start challenges, can be seamlessly integrated. Costs are determined based on the official pricing. Results are averaged over 10 seeds with a 95% confidence interval; for more setting details, refer to Section E.1. The maximum number $N$ is set uniformly to 4, according to the size of the LLM set. Budget threshold $\rho$ is 0.45 for AWC, 0.5 for SUC, and 0.3 for AIC, according to reward models and official LLM pricing. More experiments (e.g., varying budget threshold, impacts of maximum number, and comparison to scenarios without relaxation or where only two LLMs are provided, or runtime) are available in Appendix E.
Comparison Benchmarks. Comparisons include consistently utilizing the expensive ChatGPT-4 [52] and the cheap ChatGLM2 [52], the online CMAB algorithm CUCB, [67, 42], which ignores constraints, Thompson Sampling, a mainstream probabilistic Bayesian online learning approach, [39] and $\epsilon$-Greedy [2], which alternates between using empirical means and selecting uniformly based on the adaptive $\epsilon_{t}=\text{min}\{1,\frac{2\sqrt{K}}{\sqrt{t}}\}$ (Here, works of [59, 64] are not compared due to being restricted to fixed costs or linear reward). The robustness of C2MAB-V is validated by varying the parameters $\alpha_{\mu},\alpha_{c}$ with values of (0.3, 0.05), (1, 0.05), (0.3, 0.01), and (1, 0.01), referred to as (a), (b), (c), (d). Although the setting in [18, 13, 47, 73] totally differs from our online learning setting, we also pre-learned a fixed combination of multi-LLMs offline, which we applied to online queries. This enables to explore how online learning based on feedback can complement the adjustment of LLM selections in the offline domain. The results are provided in Appendix E.3 due to space constraints.
Performance Metric. To balance both reward and cost considerations, we assess performance using a reward/violation ratio, defined as the average per-round reward divided by the average per-round violation, i.e., $\frac{\sum_{\tau=1}^{t}r(S_{\tau},\boldsymbol{\mu})/t}{\sum_{\tau=1}^{t}V(\tau)/t}$, with higher ratios indicating superior performance. More extended experimental results with the varying budget threshold and a greater emphasis on performance, along with detailed results for individual reward and violation, are provided in Appendix E.2.
(a) AWC
(b) SUC
(c) AIC
Evaluation Results. The performance of the reward/ violation ratio is illustrated in Fig. 4. ChatGLM2 is excluded since its rewards are significantly low, below 0.18, 0.10, and 0.0001 in the AWC, SUC, AIC models, despite no violations. In the AWC model (Fig. 4(a)), C2MAB-V consistently outperforms other algorithms in all four parameter settings, achieving the highest reward/violation ratio. This underscores C2MAB-V’s superior ability to balance rewards against violations. Our algorithms converge in fewer than 1,000 rounds, significantly faster than the $\epsilon$-greedy algorithm, which requires nearly 10,000 rounds, even without incorporating offline information and relying entirely on online learning from scratch. The patterns observed for the SUC and AIC models (Fig. 4(b) and Fig. 4(c)) mirror those of the AWC model, further showing the robustness of C2MAB-V across different settings. When randomly evaluating C2MAB-V(c), it improves by at least 64.72% over $\epsilon$-Greedy, performs 3.9$\times$ better than Thompson Sampling, 4$\times$×that of CUCB, and 6$\times$ that of Always using ChatGPT-4 across three tasks. In summary, C2MAB-V maintains the highest reward/violation ratio under different $(\alpha_{\mu},\alpha_{c})$ for the ablation study on multi-LLM tasks aligning with our theoretical analysis.
7 Conclusion
In this paper, we introduce C2MAB-V, a cost-effective combinatorial online model with a versatile reward structure, designed to efficiently select multiple LLMs based on specific task requirements while adhering to budget constraints. C2MAB-V incorporates continual online feedback, and transform the NP-hard multi-LLM selection problem into a manageable relaxed form, with specified discretization rounding schemes designed to coordinate LLMs within a local-cloud architecture. Theoretical analysis of C2MAB-V provides novel and robust guarantees for the efficacy of the framework, including sublinear regret and a rapidly decreasing violation. Moreover, empirical evaluations with nine LLMs demonstrate C2MAB-V’s capability to balance performance with cost efficiency across three distinct types of collaborative multi-LLM tasks.
Looking ahead, there are several compelling directions for future research. For example, further development could focus on enhancing privacy protection and improving communication between multiple local servers within our local-cloud architecture. Additionally, incorporating contextual combinatorial multi-armed bandit approaches could more effectively capture user preferences.
Appendix A Description Supplement
Table 2: Summary of the main symbols.
Symbol Description
$\mathcal{K}$ Set of all LLMs, where $k$ indexes an LLM
$\mathcal{T}$ Set of discrete time intervals or rounds
$S_{t}$ Set of LLMs selected at round $t$, i.e., an action
$\mathcal{S}$ Set of all possible combinations of LLM actions
$N$ Maximum number of LLMs active simultaneously
$X_{t,k}$ Reward associated with LLM $k$ at round $t$
$y_{t,k}$ Cost associated with LLM $k$ at round $t$
$\mu_{k}$ Unknown expected performance reward for LLM $k$
$c_{k}$ Unknown expected cost for LLM $k$
$D_{q}$ Distribution over queries
$l^{\text{in}}_{k}(q)$ Number of input tokens by LLM $k$ for query $q$
$l^{\text{out}}_{k}(q)$ Random number of output tokens by LLM $k$ for query $q$
$C_{k}$ Cost per token for LLM $k$
$D^{\text{out}}_{k}(q)$ Distribution of output tokens for LLM $k$ given query $q$
$\mathcal{Q}$ Set of all queries
$\mathcal{F}_{t}$ Subset of LLMs from which feedback is observed at round $t$
$F_{t}$ Cardinality of the subset $\mathcal{F}_{t}$
$V(T)$ Total budget violation over $T$ rounds
$R(T)$ Total regret over $T$ rounds
$\rho$ Budget guarantee threshold
$\delta$ Probability parameter
$\hat{\mu}_{t,k}$ Empirical estimate of the mean reward for LLM $k$ at round $t$
$\hat{c}_{t,k}$ Empirical estimate of the cost for LLM $k$ at round $t$
$\alpha_{\mu}$ Control parameter for adjusting the reward confidence bound
$\alpha_{c}$ Control parameter for adjusting the cost confidence bound
$\rho_{t,\mu_{k}}$ Confidence radius for the reward of LLM $k$ at round $t$
$\rho_{t,c_{k}}$ Confidence radius for the cost of LLM $k$ at round $t$
$T_{t,\mu_{k}}$ Count of selections of LLM $k$ concerning the reward up to round $t$
$T_{t,c_{k}}$ Count of selections of LLM $k$ concerning the cost up to round $t$
$\tilde{z}_{k}$ Continuous variable representing the selection status of LLM $k$
$\tilde{\boldsymbol{Z}}$ Vector of $\tilde{z}_{k}$ for all LLMs
$o_{t,S}$ Probability all LLMs within action $S$ are observed at round $t$
$r^{*}$ Maximum reward for the $L$-Lipschitz reward function
$o^{*}$ Minimum observation probability across all LLM combinations
$L$ Lipschitz constant for the reward function
Appendix B Discretization Rounding Algorithms
@@
For the AWC reward, as detailed in Algorithm 2, the scheduling cloud initiates the process by setting up a vector $\boldsymbol{v}$ and identifying $K$ sets $\{\mathcal{B}_{1},\cdots,\mathcal{B}_{K}\}$ (line 2). Following this initialization, Algorithm 2 enters an iterative phase where it fine-tunes the composition of the sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$, aiming to keep the evolving solution within acceptable bounds while also seeking to optimize the reward (lines 4-10). Subsequent steps involve refining $\mathcal{B}_{1}$ by excluding all elements found in set $\mathcal{G}$, and updating $\mathcal{A}_{i+1}$ to reflect the current state of $\mathcal{B}_{1}$ (line 11), which are designed to gradually construct the final integrated solution. Finally, Algorithm 2 designates $\mathcal{A}_{K}$ as the final action of selected LLM $S_{t}$ (line 13). The validity of the resultant solution for LLM selection is underpinned by the lemma that follows (more discussions are in Section C.1).
@@
Appendix C Constraint and Reward Analysis
C.1 Discussions on Constraint Type and Extended Task Type